# Geometric Uncertainty for Detecting and Correcting Hallucinations in LLMs

### v1

@@ Abstract
Large language models demonstrate impressive results across diverse tasks but are still known to hallucinate, generating linguistically plausible but incorrect answers to questions. Uncertainty quantification has been proposed as a strategy for hallucination detection, but no existing black-box approach provides estimates for both global and local uncertainty. The former attributes uncertainty to a batch of responses, while the latter attributes uncertainty to individual responses. Current local methods typically rely on white-box access to internal model states, whilst black-box methods only provide global uncertainty estimates. We introduce a geometric framework to address this, based on archetypal analysis of batches of responses sampled with only black-box model access. At the global level, we propose Geometric Volume, which measures the convex hull volume of archetypes derived from response embeddings. At the local level, we propose Geometric Suspicion, which ranks responses by reliability and enables hallucination reduction through preferential response selection. Unlike prior dispersion methods which yield only a single global score, our approach provides semantic boundary points which have utility for attributing reliability to individual responses. Experiments show that our framework performs comparably to or better than prior methods on short form question-answering datasets, and achieves superior results on medical datasets where hallucinations carry particularly critical risks. We also provide theoretical justification by proving a link between convex hull volume and entropy.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance across diverse natural language processing tasks (Guo et al., 2025; Anthropic, 2025; Gemini Team, Google DeepMind, 2025; OpenAI, 2025) and are increasingly applied in areas such as medical diagnosis, law, and financial advice (Yang et al., 2025; Chen et al., 2024; Kong et al., 2024). Hallucinations, however, where models generate plausible but false or fabricated content, pose significant risks for adoption in high-stakes applications (Farquhar et al., 2024). Recent work, for example, finds GPT-4 hallucinating in 28.6% of reference generation tasks (Chelli et al., 2024).
@@
Uncertainty quantification (UQ) methods have been proposed to detect when models are producing unreliable outputs (Liu et al., 2025; Xiong et al., 2024). Effective UQ can serve as a critical layer of security and transparency, enabling systems to flag problematic responses and helping users exercise caution when reliability is compromised (Farquhar et al., 2024). This capability is particularly crucial for applications such as healthcare and legal services where incorrect information can cause significant harm (Huang et al., 2025; Asgari et al., 2025; Latif, 2025).
@@
UQ methods can generally be divided into two groups: sampling-based and hidden-state-based. Sampling-based methods generate multiple responses for a prompt and estimate uncertainty over the resulting batch, requiring only black-box access. Hidden-state-based methods use intermediate activations to estimate uncertainty at the single-response level, which requires white-box access. Previous work also distinguishes between external and internal uncertainty, the former arising from ambiguous queries and the latter from insufficient model knowledge (Li et al., 2025).
@@
We additionally distinguish between uncertainty estimates at the batch level, which we term global, and those at the individual response level, which we term local. Prior work has referred to the latter as ‘confidence’ (Lin et al., 2023). A comprehensive UQ method should provide estimates of both types. In high-stakes settings, local UQ enables the selection of the most reliable response from several alternatives, thereby reducing the risk of harmful hallucinations. We argue that archetypal analysis uniquely enables this unification. The extremal archetypes define the global semantic boundaries and also serve as anchors for local response-level diagnostics, something no prior convex-hull or entropy-based method provides.
@@
We introduce a geometric framework to quantify global and local uncertainty using only black-box access. Our approach is sampling-based: for a given prompt we generate a batch of responses at non-zero temperature and embed them with a sentence encoder. We then apply Archetypal Analysis (AA) (Cutler & Breiman, 1994) to identify a set of archetypes that span the embedding space. Our first contribution is Geometric Volume, a global uncertainty metric defined as the convex hull volume of the archetypes, which reflects the semantic spread of the batch.
@@
Our second contribution is Geometric Suspicion, a local uncertainty measure that accounts for the distribution of responses within the spanned volume and how each response is reconstructed from the archetypes. To our knowledge, this is the first sampling-based black-box method that differentiates high and low uncertainty responses within a batch.
@@
  • We introduce Geometric Suspicion, the first black-box, sampling-based local uncertainty method, which reduces hallucinations by guiding Best-of-N response selection.
@@ 2 Related Works
Our work is closely related to Semantic Volume (Li et al., 2025), which also analyzes the semantic dispersion of a batch of natural language outputs. Specifically, Semantic Volume computes the determinant of the Gram matrix formed from a batch of embeddings, where a low value ($\log\det(V^{\top}V)$) corresponds to low uncertainty and vice versa.
@@
Several recent works use convex hull area over response embeddings as a proxy for uncertainty, projecting into two dimensions and summing hulls around clusters (Catak & Kuzlu, 2024; Catak et al., 2024). These works however vastly simplify the semantic space by projecting into only two dimensions with principal component analysis. They also first cluster responses before computing and summing the convex hull area around each cluster, which ignores how far apart in embedding space separate clusters may be. In doing so crucial information is lost regarding the variation in response meaning a model is likely to give to a certain prompt.
@@
Semantic entropy detects hallucinations by clustering semantically equivalent responses using bidirectional entailment, then computing entropy over those clusters (Farquhar et al., 2024). Self-consistency methods have built on their work and are emerging as a dominant paradigm (Taubenfeld et al., 2025; Wan et al., 2025; Savage et al., 2024), where multiple responses are sampled and their agreement is used as an uncertainty signal. Again, neither of these works offer additional methods for local UQ.
@@
Finally, numerous methods have been proposed to estimate uncertainty with white-box model access, using for instance token probabilities, logits or hidden layer activations without requiring additional sampling. These methods include minimum token probability, average token probability, and more advanced techniques that account for the semantic importance of individual tokens (Xia et al., 2025; Zhang et al., 2025; Liu et al., 2024; Malinin & Gales, 2020; Quevedo et al., 2024).
To our knowledge, we are the first to introduce a unified framework that connects global dispersion with local attribution. Unlike prior methods that collapse geometry into a single global score, our use of archetypal analysis yields interpretable anchor points that both define batch-level uncertainty and enable principled response-level attribution in a black-box setting. This enables fine-grained hallucination detection and improves interpretability without requiring access to internal model states.
@@ 3.2 Global Uncertainty via Geometric Volume
At a high level, we use the set of responses $\mathbf{X}$ to evaluate the uncertainty associated with $r_{\text{default}}$. Our global uncertainty metric is based on archetypal analysis (AA) (Cutler & Breiman, 1994), an unsupervised method for finding representative archetypes that lie on the convex hull of the data. Unlike clustering methods that identify centroids within the data cloud, archetypes represent extremal points, thereby capturing the semantic boundaries of possible model responses.
@@
This bi-convex constraint leads to the following optimization problem (Abrol & Sharma, 2020):
@@
In practice, Eq. 1 is solved iteratively via block-coordinate descent (Chen et al., 2014), alternating between updates of $\mathbf{A}$ and $\mathbf{B}$ until convergence. This procedure ensures that the resulting archetypes capture extremal directions of the data, approximating the vertices of the convex hull.
@@
3.3 Local Uncertainty Detection
This global score $H_{G}(\mathbf{X})$ allows us to classify hallucinations successfully, but only facilitates deferral as a method of reducing hallucinations. This is because while the global score indicates if a response set is uncertain, it does not attribute uncertainty measures to individual responses within the set. Sensible measures of local uncertainty could allow reduction in hallucination rates by preferential selection of more certain responses, i.e. those less likely to be hallucinations, from the response set. To address this we introduce a simple local uncertainty measure, which we term Geometric Suspicion, composed of three separate terms derived from the archetypal coefficient matrix $\mathbf{A}$ and the answer embeddings $\mathbf{X}$.
To derive and support our choice of terms, we conduct an analysis of the datasets generated for our global uncertainty experiments, which can be found in Appendix Section D.
@@ 3.3 Local Uncertainty Detection
This metric operates directly on the set of response embeddings $\mathbf{X}$. For each embedding $\mathbf{x}_{i}\in\mathbf{X}$, we identify its $k$-nearest neighbours. The score is the average Euclidean distance to these neighbours:
$L(r_{i})=\frac{1}{k}\sum_{\mathbf{x}_{j}\in\mathcal{N}_{k}(\mathbf{x}_{i})}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}.$ (3)
Here, $\mathcal{N}_{k}(\mathbf{x}_{i})$ is the set of the $k$-nearest neighbours to embedding $\mathbf{x}_{i}$. The intuition is that correct, coherent responses will form dense clusters (resulting in a low distance score), while hallucinations or semantic outliers will be isolated in sparse regions, yielding a high score.
@@
While Local Density is effective at finding isolated points, it can fail if a model produces a coherent but incorrect cluster of responses. To address this, our second metric establishes a global geometric context. It measures how far a response embedding $\mathbf{x}_{i}$ is from the semantic consensus point $\mathbf{x}_{c}$, defined as the average embedding across the response set:
$D(r_{i})=\|\mathbf{x}_{i}-\mathbf{x}_{c}\|_{2},\quad\text{where}\quad\mathbf{x}_{c}=\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{j}.$ (4)
This score identifies responses that are semantically distant from the central theme of the entire answer set. A high score signifies that a response, even if part of a dense local cluster, belongs to a region that is globally anomalous.
@@
The final metric leverages Archetypal Analysis to identify responses that rely on fringe boundaries of the semantic space. The $i$-th row of $\mathbf{A}$, denoted $\bm{\alpha}_{i}=[A_{i1},\dots,A_{iK}]$ contains coefficients describing how the response embedding $\mathbf{x}_{i}$ is reconstructed from the set of learned archetypes $\mathbf{Z}$. The score for a response $r_{i}$ is high if its reconstruction coefficients $\mathbf{A}_{ik}$ depend heavily on archetypes that are, on average, rarely used to explain the rest of the data.:
@@
Final Suspicion Score
The final ‘suspicion’ score, $S(r_{i})$, is created by combining the evidence from the three individual terms in a non-parametric fashion. For each of the three metrics, we rank all responses in the batch from least suspicious (Rank 1) to most suspicious (Rank $n$). The final suspicion score for a response is the sum of its ranks across the three metrics:
$S(r_{i})=\text{rank}_{L}(r_{i})+\text{rank}_{D}(r_{i})+\text{rank}_{U}(r_{i}),$ (6)
where $\text{rank}_{M}(r_{i})$ is the rank of response $r_{i}$ according to metric $M$ within the batch. By summing the ranks, we create a robust score that is insensitive to the individual scales of the metrics and identifies responses that are consistently flagged as anomalous. We then select the response with the minimum combined rank as the most plausible answer.
@@ 3.4 Hallucination Detection
Our geometric volume metric can be used to detect hallucinations at the response set level. Given a set of responses with embedding matrix $\mathbf{X}$, we classify the entire set as hallucinated if the global uncertainty exceeds a threshold $\tau$, i.e., if $H_{G}(\mathbf{X})>\tau$. The threshold $\tau$ is a tunable hyperparameter, which can be optimized on a small labeled validation set to maximize a classification objective such as F1-score.
3.5 Reducing Hallucination Rates with Local Uncertainty
@@
To quantify the efficacy of this selection strategy, we measure the net reduction in the hallucination rate across a dataset. First, we define the baseline hallucination rate, $H_{\text{baseline}}$, as the proportion of questions in a test set for which the model’s default answer $r_{\text{default}}$ was a hallucination. Next, we define the post-selection hallucination rate, $H_{\text{BoN-U}}$, as the proportion of questions where the answer $r^{*}$ selected by our uncertainty-guided method is a hallucination.
@@ 4.1 Benchmarks and Baselines
We evaluate on five benchmarks covering both external and internal uncertainty: CLAMBER (ambiguous prompts), TriviaQA (short-form QA), ScienceQA (scientific reasoning), MedicalQA (high-stakes medical QA), and K-QA (real-world medical scenarios). These benchmarks span ambiguous queries, factual QA, and long-form medical domains. Dataset construction details are provided in Section C of the Appendix. We compare our method with Semantic Volume (Li et al., 2025), Semantic Entropy (Farquhar et al., 2024), and P(true), which estimates the probability that a model’s generated response is correct by directly asking the model itself to self-assess its confidence (Lin et al., 2023).
@@
Local Uncertainty Evaluation
We evaluate our local uncertainty quantification methods on the same datasets curated for the global Internal Uncertainty experiments. We first filter each dataset to only contain questions where the set of sampled responses contains examples of both hallucinations and non-hallucinations. For each sample in the perturbed responses, we then compute the local uncertainty score using the archetypal reconstruction matrix $\mathbf{A}$ and embeddings $\mathbf{X}$.
@@
All Valid ($0<H<1$) Mid-Valid ($0.33<H<0.67$)
Dataset Model $\DeltaHR Baseline HR BoN HR Δ\DeltaHR$
K-QA
GPT-3.5-Turbo 60.41.2 29.15.3 31.34.4 65.08.7 40.310.6 24.74.3
GPT-4o Mini 48.52.2 33.42.2 15.13.4 54.52.5 36.97.4 17.77.0
Llama3.1-8b 66.21.2 51.51.8 14.73.0 62.40.9 45.94.7 16.55.3
Qwen3-8b 55.52.9 36.84.6 18.77.3 64.82.5 35.22.5 29.64.9
MedicalQA
GPT-3.5-Turbo 52.51.7 43.71.1 8.90.7 57.42.3 39.52.4 17.94.5
GPT-4o Mini 48.71.5 39.70.8 9.01.0 50.93.8 42.03.3 8.90.6
Llama3.1-8b 52.91.8 47.22.3 5.70.8 42.93.3 29.45.5 13.52.3
Qwen3-8b 52.03.2 44.62.1 7.41.4 50.24.3 37.42.9 12.86.0
ScienceQA
GPT-3.5-Turbo 80.64.7 77.87.0 2.82.5 77.94.8 68.77.9 9.25.2
GPT-4o Mini 91.24.9 85.05.3 6.21.0 96.53.0 76.65.0 19.92.0
Llama3.1-8b 85.20.3 77.43.3 7.83.2 85.82.6 65.59.7 20.38.5
Qwen3-8b 92.31.8 88.61.6 3.73.3 100.00.0 80.818.8 19.218.8
TriviaQA
GPT-3.5-Turbo 46.20.8 40.20.4 6.00.7 47.94.2 26.92.7 21.14.3
GPT-4o Mini 52.91.0 46.20.5 6.61.5 44.32.2 22.80.5 21.51.7
Llama3.1-8b 55.71.5 53.71.3 2.01.7 47.92.5 38.52.3 9.41.1
Qwen3-8b 45.00.2 45.60.9 -0.50.7 30.10.9 27.41.5 2.71.8

### v2

@@ Abstract
Large language models demonstrate impressive results across diverse tasks but are still known to hallucinate, generating linguistically plausible but incorrect answers to questions. Uncertainty quantification has been proposed as a strategy for hallucination detection, requiring estimates for both global uncertainty (attributed to a batch of responses) and local uncertainty (attributed to individual responses). While recent black-box approaches have shown some success, they often rely on disjoint heuristics or graph-theoretic approximations that lack a unified geometric interpretation. We introduce a geometric framework to address this, based on archetypal analysis of batches of responses sampled with only black-box model access. At the global level, we propose Geometric Volume, which measures the convex hull volume of archetypes derived from response embeddings. At the local level, we propose Geometric Suspicion, which leverages the spatial relationship between responses and these archetypes to rank reliability, enabling hallucination reduction through preferential response selection. Unlike prior methods that rely on discrete pairwise comparisons, our approach provides continuous semantic boundary points which have utility for attributing reliability to individual responses. Experiments show that our framework performs comparably to or better than prior methods on short form question-answering datasets, and achieves superior results on medical datasets where hallucinations carry particularly critical risks. We also provide theoretical justification by proving a link between convex hull volume and entropy.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance across diverse natural language processing tasks (guo2025deepseek; anthropic_claude4_2025; deepmind_gemini2.5_report_2025; openai_o3_o4mini_system_card_2025) and are increasingly applied in areas such as medical diagnosis, law, and financial advice (yang2025application; chen2024survey; kong2024large). Hallucinations, however, where models generate plausible but false or fabricated content, pose significant risks for adoption in high-stakes applications (farquhar2024detecting). Recent work, for example, finds GPT-4 hallucinating in 28.6% of reference generation tasks (chelli2024hallucination).
@@
Uncertainty quantification (UQ) methods have been proposed to detect when models are producing unreliable outputs (liu2025uncertainty; xiong2024efficient). Effective UQ can serve as a critical layer of security and transparency, enabling systems to flag problematic responses and helping users exercise caution when reliability is compromised (farquhar2024detecting). This capability is particularly crucial for applications such as healthcare and legal services where incorrect information can cause significant harm (huang2025survey; asgari2025framework; latif2025hallucinations).
@@
UQ methods can generally be divided into two groups: sampling-based and hidden-state-based. Sampling-based methods generate multiple responses for a prompt and estimate uncertainty over the resulting batch, often requiring only black-box access. Hidden-state-based methods use intermediate activations to estimate uncertainty at the single-response level, which requires white-box access. Previous work also distinguishes between external and internal uncertainty, the former arising from ambiguous queries and the latter from insufficient model knowledge (li2025semantic).
@@
We additionally distinguish between uncertainty estimates at the batch level, which we term global, and those at the individual response level, which we term local. Prior work has referred to the latter as ‘confidence’ (lin2024generatingconfidenceuncertaintyquantification). A comprehensive UQ method should provide estimates of both types. In high-stakes settings, local UQ enables the selection of the most reliable response from several alternatives, thereby reducing the risk of harmful hallucinations. We argue that archetypal analysis uniquely enables this unification. The extremal archetypes define the global semantic boundaries and also serve as anchors for local response-level diagnostics, something no prior convex-hull or entropy-based method provides.
@@
We introduce a geometric framework to quantify global and local uncertainty using only black-box access. Our approach is sampling-based: for a given prompt we generate a batch of responses at non-zero temperature and embed them with a sentence encoder. We then apply Archetypal Analysis (AA) (cutler1994archetypal) to identify a set of archetypes that span the embedding space. Our first contribution is Geometric Volume, a global uncertainty metric defined as the convex hull volume of the archetypes, which reflects the semantic spread of the batch.
@@
Our second contribution is Geometric Suspicion, a local uncertainty measure that accounts for the distribution of responses within the spanned volume and how each response is reconstructed from the archetypes. By deriving local suspicion directly from the boundary conditions of the global volume, our method offers a geometrically grounded alternative to heuristic graph measures for differentiating high and low uncertainty responses.
@@
  • We introduce Geometric Suspicion, a black-box, sampling-based local uncertainty method integrated within the same geometric framework, which reduces hallucinations by guiding Best-of-N response selection.
@@ 2 Related Works
Our work is closely related to Semantic Volume (li2025semantic), which also analyzes the semantic dispersion of a batch of natural language outputs. Specifically, Semantic Volume computes the determinant of the Gram matrix formed from a batch of embeddings, where a low value ($\log\det(V^{\top}V)$) corresponds to low uncertainty and vice versa.
@@
Several recent works use convex hull area over response embeddings as a proxy for uncertainty, projecting into two dimensions and summing hulls around clusters (catak2024uncertainty; catak2024improving). These works however vastly simplify the semantic space by projecting into only two dimensions with principal component analysis. They also first cluster responses before computing and summing the convex hull area around each cluster, which ignores how far apart in embedding space separate clusters may be. In doing so crucial information is lost regarding the variation in response meaning a model is likely to give to a certain prompt.
@@
Semantic entropy detects hallucinations by clustering semantically equivalent responses using bidirectional entailment, then computing entropy over those clusters (farquhar2024detecting). Self-consistency methods have built on their work and are emerging as a dominant paradigm (taubenfeld2025confidence; wan2025reasoning; savage2024large), where multiple responses are sampled and their agreement is used as an uncertainty signal. Again, neither of these works offer additional methods for local UQ.
Semantic Space Confidence:
Several approaches have been proposed to assess the local uncertainty, or confidence, of individual answers by analysis of the semantic answer embeddings. Semantic Density (qiu2024semanticdensityuncertaintyquantification) approximates the probability distribution over answers in semantic space and assigns high confidence to answers lying in high probability regions, but uses model likelihoods to do so and as such is not a black-box method. lin2024generatingconfidenceuncertaintyquantification propose several black-box methods to assess global and local uncertainty, including Degree and Eccentricity. These methods use graphs computed from pairwise comparisons between answer representations, and as such lose the full semantic context of the response set, which is captured by our method.
@@
Finally, numerous methods have been proposed to estimate uncertainty with white-box model access, using for instance token probabilities, logits or hidden layer activations without requiring additional sampling. These methods include minimum token probability, average token probability, and more advanced techniques that account for the semantic importance of individual tokens (xia2025survey; zhang2025token; liu2024uncertainty; malinin2020uncertainty; quevedo2024detecting).
Unlike prior methods that collapse geometry into a single global score, our use of archetypal analysis yields interpretable anchor points that both define batch-level uncertainty and enable principled response-level attribution in a black-box setting. This enables fine-grained hallucination detection and improves interpretability without requiring access to internal model states.
@@ 3.2 Global Uncertainty via Geometric Volume
At a high level, we use the set of responses $\mathbf{X}$ to evaluate the uncertainty associated with $r_{\text{default}}$. Our global uncertainty metric is based on archetypal analysis (AA) (cutler1994archetypal), an unsupervised method for finding representative archetypes that lie on the convex hull of the data. Unlike clustering methods that identify centroids within the data cloud, archetypes represent extremal points, thereby capturing the semantic boundaries of possible model responses.
@@
This bi-convex constraint leads to the following optimization problem (icml_aa):
@@
In practice, Eq. 1 is solved iteratively via block-coordinate descent (CVPR_AA), alternating between updates of $\mathbf{A}$ and $\mathbf{B}$ until convergence. This procedure ensures that the resulting archetypes capture extremal directions of the data, approximating the vertices of the convex hull.
@@
3.3 Simplicial Volume as a Proxy for Uncertainty
We now formalize this relationship with Theorem 1, which states that the volume of the archetypal simplex provides an upper bound on the differential entropy of any distribution supported within it, thereby linking geometric dispersion to semantic uncertainty. We provide the proof in the Appendix.
Theorem 1.
Proof.
See Appendix  A.1. ∎
3.4 Local Uncertainty via Geometric Suspicion
@@
The global score $H_{G}(\mathbf{X})$ allows us to classify hallucinations successfully but only facilitates deferral as a method of reducing hallucinations, as it does not attribute uncertainty to individual responses within a set. Sensible local scores would enable hallucination reduction by preferentially selecting more certain responses from the response set. While recent graph-based approaches have demonstrated the utility of local metrics (lin2024generatingconfidenceuncertaintyquantification), we propose Geometric Suspicion, a local measure explicitly derived from the convex hull geometry used for our global metric. This is built from three per–response terms derived from the archetypal coefficients $\mathbf{A}$ and the answer embeddings $\mathbf{X}$.
@@
For each response embedding $\mathbf{x}_{i}\in\mathbf{X}$, let $\mathcal{N}_{k}(\mathbf{x}_{i})$ be its $k$-nearest neighbours. We score
$L(r_{i})=\frac{1}{k}\sum_{\mathbf{x}_{j}\in\mathcal{N}_{k}(\mathbf{x}_{i})}\big\|\mathbf{x}_{i}-\mathbf{x}_{j}\big\|_{2},$ (4)
so that isolated (locally sparse) responses receive larger values.
@@
To place each response in a global geometric context, we measure the distance to the batch consensus
$D(r_{i})=\big\|\mathbf{x}_{i}-\mathbf{x}_{c}\big\|_{2},\qquad\mathbf{x}_{c}=\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{j}.$ (5)
Large values indicate semantic deviation from the central tendency of the answer set.
@@
The $i$-th row of $\mathbf{A}$, denoted $\bm{\alpha}_{i}=[A_{i1},\dots,A_{iK}]$ contains coefficients describing how the response embedding $\mathbf{x}_{i}$ is reconstructed from the set of learned archetypes $\mathbf{Z}$. The score for a response $r_{i}$ is high if its reconstruction coefficients $\mathbf{A}_{ik}$ depend heavily on archetypes that are, on average, rarely used to explain the rest of the data.:
@@
Final Suspicion Score (Fisher combination)
We convert each per–metric score into a within–batch empirical right–tail pp–value. For metric $M\in\{L,D,U\}$ and response $r_{i}$,
$p_{i}^{(M)}\;=\;\frac{1+\left|\,\{\,j\in\{1,\dots,n\}:\;s_{j}^{(M)}\geq s_{i}^{(M)}\,\}\,\right|}{\,n+1\,}$, (7)
where $s_{i}^{(M)}$ is oriented so that larger values indicate greater suspicion. We then combine the three pp–values using Fisher’s method to obtain the final suspicion statistic
$S(r_{i})\;=\;-2\sum_{M\in\{L,D,U\}}\log p_{i}^{(M)}.$ (8)
Relation to prior local UQ baselines and intuition
Our two geometric terms align closely with established graph-based baselines, namely Degree and Eccentricity (lin2024generatingconfidenceuncertaintyquantification). Local Density is a pointwise analogue of Degree on an answer–answer affinity graph: degree is a (monotone) transform of local similarity mass, while equation 4 measures the corresponding local spacing in $\mathbf{X}$. Distance from Consensus mirrors Eccentricity (distance to a central embedding in a low-dimensional spectral space): both penalize globally off-centre responses, even when they form dense clusters. Usage Rarity contributes complementary geometric signal from archetypal analysis by highlighting reliance on rarely used (hull-boundary) archetypes, which our global theory links to increased dispersion. Together, the three terms provide local sparsity (degree–like), global deviation (eccentricity–like), and boundary reliance (archetypal) evidence.
To derive and support our specific choice of terms, we conduct an analysis of the datasets generated for our global uncertainty experiments, which can be found in Appendix Section D. In Figure 3, we show examples where component metrics used in isolation would not select an optimal response from a batch, whereas their combination overcomes individual limitations. Degree / Local Density used in isolation, for instance, struggle in situations where there exist multiple dense clusters in the semantic response set.
@@ 3.5 Hallucination Detection
Our geometric volume metric can be used to detect hallucinations at the response set level. Given $r_{\text{default}}$ and a corresponding set of responses with embedding matrix $\mathbf{X}$, we classify $r_{\text{default}}$ as hallucinated if the global uncertainty exceeds a threshold $\tau$, i.e., if $H_{G}(\mathbf{X})>\tau$. The threshold $\tau$ is a tunable hyperparameter, which can be optimized on a small labeled validation set to maximize a classification objective such as F1-score.
3.6 Hallucination Correction
@@
To quantify the efficacy of this selection strategy, we measure the net reduction in the hallucination rate across a dataset. First, we define the baseline hallucination rate, $H_{\text{baseline}}$, as the proportion of questions in a test set for which the model’s default answer $r_{\text{default}}$ was a hallucination.
Next, we define the post-selection hallucination rate, $H_{\text{BoN-U}}$, as the proportion of questions where the answer $r^{*}$ selected by our uncertainty-guided method is a hallucination.
@@ 4.1 Benchmarks and Baselines
We evaluate on five benchmarks covering both external and internal uncertainty: CLAMBER (ambiguous prompts), TriviaQA (short-form QA), ScienceQA (scientific reasoning), MedicalQA (high-stakes medical QA), and K-QA (real-world medical scenarios). These benchmarks span ambiguous queries, factual QA, and long-form medical domains. Dataset construction details are provided in Section B of the Appendix. We compare our method with Semantic Volume (li2025semantic), Semantic Entropy (farquhar2024detecting), and P(true), which estimates the probability that a model’s generated response is correct by directly asking the model itself to self-assess its confidence (lin2024generatingconfidenceuncertaintyquantification).
@@
We evaluate our local uncertainty quantification methods on the same datasets curated for the global Internal Uncertainty experiments. We first filter each dataset to only contain questions where the set of sampled responses contains examples of both hallucinations and non-hallucinations. For each sample in the perturbed responses, we then compute the local uncertainty score using the archetypal reconstruction matrix $\mathbf{A}$ and embeddings $\mathbf{X}$. We compare our method with Degree and Eccentricity (lin2024generatingconfidenceuncertaintyquantification), which are based on graphs derived from pairwise response comparisons. We report $\Delta H$ and AUARC (Area Under the Accuracy–Rejection Curve). The latter is computed per question by ranking responses by uncertainty, progressively rejecting the top-$k$, measuring accuracy on the retained set, and averaging over all rejection rates ($k=0,\dots,n-1$). A random predictor yields AUARC equal to the base accuracy.
@@
Method TriviaQA ScienceQA MedicalQA K-QA
[…]
GPT-4o Mini
Eccentricity -1.01.1 48.31.0 4.90.9 16.44.9 2.71.0 55.72.1 4.94.0 58.50.2
Degree 7.72.5 51.90.3 5.70.6 16.84.6 7.03.7 57.82.4 15.11.9 63.70.3
Ours 6.91.1 51.90.2 6.21.0 17.34.8 8.91.6 58.61.8 15.32.9 64.10.2
GPT-3.5-Turbo
Eccentricity 1.02.0 55.80.2 5.51.4 25.85.7 1.72.0 51.51.2 20.84.5 64.30.2
Degree 5.70.8 58.11.0 4.52.9 25.16.6 8.71.4 54.40.6 32.23.2 68.81.6
Ours 6.00.6 58.10.7 2.82.5 24.56.8 8.80.6 55.00.5 31.33.0 68.31.7
Qwen3-8b
Eccentricity -9.21.8 47.30.9 2.22.3 13.20.9 -1.01.6 49.72.9 9.84.8 55.22.8
Degree -0.71.6 51.60.8 3.73.3 13.70.1 7.52.2 53.42.3 15.38.6 58.53.6
Ours -0.50.7 51.60.8 3.73.3 14.80.5 8.01.1 54.12.4 19.07.2 59.93.2
Llama3.1-8b
Eccentricity -2.60.3 42.20.2 8.21.8 25.01.6 -7.52.9 41.80.8 9.23.9 44.80.6
Degree 4.52.1 45.50.8 7.03.0 24.52.1 4.03.0 47.20.8 12.33.3 46.41.6
Ours 2.11.7 44.70.6 8.02.9 25.52.0 5.21.2 48.61.3 15.33.6 46.91.3